# A Survey on Diffusion Language Models

### v1

@@ 6 Performance Study
Our findings suggest that DLMs generally perform competitively with AR models of comparable size. On general language understanding benchmarks such as PIQA and HellaSwag, models like LLaDA achieve performance that is slightly below or on par with AR models such as LLaMA2 [4] and Qwen2.5 [166]. However, DLMs exhibit stronger performance in math and science-related benchmarks, including GSM8K, GPQA [167], and MATH [168], where models such as LLaDA and Dream consistently outperform similarly sized AR counterparts. In multimodal tasks, models like MMaDA [31] and LLaDA-V [29] often surpass AR-based multimodal models, highlighting the potential of DLMs in unified and cross-modal reasoning. On code generation tasks, DLMs also demonstrate competitive capabilities. Notably, DiffuCoder [84] achieves competitive HumanEval performance among open-source models, illustrating the potential of DLMs in structured, logic-heavy domains. Furthermore, closed-source DLMs such as Gemini Diffusion [33] and Mercury [32] achieve state-of-the-art results among all DLMs, rivalling top-tier AR models like GPT-4o.
@@ 7.2 Code Generation
Although DLMs are rarely explicitly designed for code generation, the global planning and iterative refinement capabilities of them are particularly well-suited for the non-sequential nature of code generation. Foundational models like DiffuCoder [84], a 7B open-source model, have been developed specifically for this domain. DiffuCoder’s analysis reveals unique decoding behaviors, such as generation order becoming more flexible at higher temperatures. It also proposes coupled-GRPO, a novel sampling scheme that constructs complementary mask noise for completions used in training, which significantly improves the model’s performance on code generation tasks. Building on the reasoning aspect, DCoLT [82] treats the entire reverse diffusion process as a form of non-linear, ”lateral” thinking. With outcome-based RL and unmasking policy module, it achieves strong results on complex coding tasks. Dilated Unmasking Scheduler (DUS) [119] offers an inference-only, planner-free method that unmasks tokens in a non-adjacent pattern to minimize an upper bound on joint entropy gain at each denoising step, achieving promising results on code generation while improving speed-quality trade-off. Demonstrating the real-world potential of DLMs’ speed, Mercury Coder [32] is a commercial-scale diffusion model that achieves state-of-the-art throughput, outperforming speed-optimized autoregressive models by up to 10 times while maintaining comparable quality on major code benchmarks.
@@ 8.1 Major Challenges
3) Long Sequence and Dynamic-Length Generation. DLMs are typically trained to denoise fixed-length sequences under a diffusion-based objective, which makes it challenging to generalize to longer or dynamically sized sequences at inference time. Most existing DLMs are limited to a maximum context length of 4,096 tokens, and widely used extrapolation techniques in AR models for longer sequences remain underexplored in the DLM setting. This limitation hinders the applicability of DLMs in tasks requiring long-context understanding or complex reasoning. In addition, DLMs generally require the generation length to be predetermined during inference, making them ill-suited for dynamic-length generation. Although DLMs can predict an [EOS] token and omit displaying tokens generated afterward, the entire sequence is still fully updated throughout the denoising process, regardless of whether the generation has logically ended, which leads to unnecessary computational overhead. In addition, masked DLMs utilize full bidirectional attention at every denoising step, which incurs a computational cost of $\mathcal{O}(N^{2})$ per step, where $N$ is the sequence length. Assuming a fixed number of tokens are unmasked at each step, the total number of denoising steps scales linearly with $N$, leading to an overall inference complexity of $\mathcal{O}(N^{3})$. Without architectural optimizations such as KV-Cache, this cubic time complexity severely limits the scalability of DLMs for long-sequence generation in real-world applications.

### v2

@@ 6 Performance Study
Our findings suggest that DLMs generally perform competitively with AR models of comparable size. On general language understanding benchmarks such as PIQA and HellaSwag, models like LLaDA achieve performance that is slightly below or on par with AR models such as LLaMA2 [4] and Qwen2.5 [194]. However, DLMs exhibit stronger performance in math and science-related benchmarks, including GSM8K, GPQA [195], and MATH [196], where models such as LLaDA and Dream consistently outperform similarly sized AR counterparts. In multimodal tasks, models like MMaDA [31] and LLaDA-V [29] often surpass AR-based multimodal models, highlighting the potential of DLMs in unified and cross-modal reasoning. On code generation tasks, DLMs also demonstrate competitive capabilities. Notably, DiffuCoder [89] achieves competitive HumanEval performance among open-source models, illustrating the potential of DLMs in structured, logic-heavy domains. Furthermore, closed-source DLMs such as Gemini Diffusion [33] and Mercury [32] achieve state-of-the-art results among all DLMs, rivaling top-tier AR models like GPT-4o.
@@
Recent scaling studies further show that DLMs tend to outperform AR models in data-constrained, multi-epoch regimes, likely because their any-order denoising objective enables more effective reuse of limited data [171].
@@ 7.2 Code Generation
[…]
@@ 7.3 Biological and Scientific Applications
7.4 Robotics
Recently, DLM-based vision-language-action (VLA) models have demonstrated strong potential in unifying perception, reasoning, and control. Built upon LLaDA, LLaDA-VLA [147] incorporates localized special-token classification alongside hierarchical action-structured decoding, leading to significant improvements over autoregressive VLA baselines in both simulation and real-world evaluations. dVLA [147] leverages pretrained MMaDA as a diffusion backbone to jointly generate visual subgoal images, textual chain-of-thought, and discretized actions, further introducing prefix-attention masking and KV caching for efficient long-horizon manipulation. Unified Diffusion VLA (UD-VLA) [148] proposes a Joint Discrete Denoising Diffusion Process that synchronously denoises future image and action tokens within a shared token space, achieving state-of-the-art performance on benchmarks with substantially improved inference speed.
@@ 8.1 Major Challenges
3) Long Sequence and Dynamic-Length Generation. DLMs are typically trained to denoise fixed-length sequences under a diffusion-based objective, which makes it challenging to generalize to longer or dynamically sized sequences at inference time. Most existing DLMs are limited to a maximum context length of 4,096 tokens, and widely used extrapolation techniques in AR models for longer sequences remain underexplored in the DLM setting. This limitation hinders the applicability of DLMs in tasks requiring long-context understanding or complex reasoning. In addition, DLMs generally require the generation length to be predetermined during inference, making them ill-suited for dynamic-length generation. Although DLMs can predict an [EOS] token and omit displaying tokens generated afterward, the entire sequence is still fully updated throughout the denoising process, regardless of whether the generation has logically ended, which leads to unnecessary computational overhead. Recent work has begun to address this limitation through both training-based [201, 202] and training-free [203, 204] approaches. In addition, masked DLMs utilize full bidirectional attention at every denoising step, which incurs a computational cost of $\mathcal{O}(N^{2})$ per step, where $N$ is the sequence length. Assuming a fixed number of tokens are unmasked at each step, the total number of denoising steps scales linearly with $N$, leading to an overall inference complexity of $\mathcal{O}(N^{3})$. Without architectural optimizations such as KV-Cache, this cubic time complexity severely limits the scalability of DLMs for long-sequence generation in real-world applications.